Consider the production function Q = 2(KL)0.5

a)     What is the marginal product of labour and capital

b)     What is the marginal rate of technical substitution of labor for capital


c)     What is the elasticity of substitution at a point K = 1, L = 1 if we increase K by one unit?


 

1
Expert's answer
2021-05-06T15:53:20-0400

With the given function of production "Q=2(KL)^{0.5}"

1) To find the products (MPK and also MPL), we use this formula. 

"Q=2(KL)^{0.5}"


"MPK=\frac{∂Q}{∂K}=L^{0.5}"


"MPK=L^{0.5}"


"MPL=\frac{\delta Q}{\delta L}=(KL)^{-0.5}"


2) To find the value of substitution (technical) of labor for capital

"MRTS=\frac {MPL}{MPK}"


"MPTS=\frac{KL^{-0.5}}{L^{0.5}}"


"MRTS=\frac {1}{KL^{0.5}}"


3) To find how much the substitution is elastic


"ε=(\frac{\Delta \frac{1}{k}}{\Delta MRTS})\frac{MRTS}{\frac{1}{k}}"


we know that "MRTS=\frac {1}{KL^{0.5}}"


taking the derivative

"MRTS=\frac {1}{KL^{0.5}}"


"MRTS=0.5(KL)^{-1.5}"


"MRTS=\frac {1}{0.5(KL)^{1.5}}"


"MRTS=\frac {1}{0.5(KL)^{1.5}}"


"\frac{\Delta \frac{1}{k}}{\Delta MRTS}=\frac {1}{0.5(KL)^{1.5}}"


"ε=\frac{1}{0.5KL}"


"ε=\frac{1}{0.5(1\times 1)}"


"\u03b5=2"


the elasticity of substitution would be 2

but, if K is raised by 1, then


"ε=\frac{1}{0.5KL}"



"ε=\frac{1}{0.5(2 \times 1)}"


"\u03b5=1"


Now the elasticity will fall to 1